If roots of equation x³+ax²+bx+c=0 (a, b, c not equal to 0) are three consecutive integers, then a(b-2)/equals to

Grade:11

1 Answers

subrat
6 years ago
You must be asking a(b-2)/c. If that`s the case then let us suppose roots of the equation to be k-1,k,k+1. Then sum of roots i.e. is 3k=-a , sum-product is 3k^2 -1 = b and product of roots is k^3 - k=c. So a(b-2)/c = -9.